OpenStudy (anonymous):

Solve each equation 180=220-k

OpenStudy (anonymous):

formula! if A and B are some numbers then if we have: B = A - k we solve for k: B+k = A and then k = A - B

OpenStudy (anonymous):

I am so confused

OpenStudy (radar):

Subtract 220 from both sides 180-220=220-220-k now consolidate terms and simplify -40=-k multiply by (both sides) by -1 40=k or k=40

OpenStudy (radar):

Please note, you want to isolate the unknown by doing algebraic operations to both sides of the equation......Must do to both sides or the equation will become invalid.

OpenStudy (radar):

Was there any operation you did not understand?

OpenStudy (radar):

I am taking you non response as an "I understand" bye

OpenStudy (anonymous):

thank u